I am trying to help my child with:
2(x+1) + 3x = 2x -4-x +16. The problem we are having is that the teacher says to solve it a different way and we got x=4. I think the teacher is using distributive property. Can you help?

Found 2 solutions by edjones, scott8148:
Answer by edjones(7569)

2(x+1) + 3x = 2x -4-x +16
2x+2+3x=2x-4-x+16
5x+2=x+12
4x=10
x=10/4=5/2
.
Ed

Answer by scott8148(6628)

distributing ___ 2x + 2 + 3x = 2x - 4 - x + 16

consolidating like terms ___ 5x + 2 = x + 12

subtracting x+2 ___ 4x = 10

dividing by 4 ___ x = 10/4 = 5/2